OpenStudy (anonymous):

1+1/3a=6 wat is the solution of the equation

OpenStudy (anonymous):

why dont you try to do it yourself instead of keep asking?

OpenStudy (anonymous):

Multiply everything by 3a firstly.

OpenStudy (anonymous):

Now let's do it step by step to see where you have any doubt

OpenStudy (anonymous):

\[a=\frac{1}{15}\] Please try working it out yourself too.

OpenStudy (anonymous):

1+(1/3)a=6 (1/3)a=6-1 (1/3)a=5 a=5*(3/1) a=15

OpenStudy (anonymous):

nobody ask you to answer. Jesus. And 1/15 isn't the answer. The computer said your wrong. The thing is if I wanted to do it I would because I can. I ask questions to get my level to go up and any one who comments my Q wrong or right I give medals so they can go a level up. Also I Often write a question and answer it my self but ask everyone if I got it right. Thank you.

OpenStudy (anonymous):

Thanks for your help Dalvorn but there is no 3a only a 1/3a

OpenStudy (anonymous):

women... lol "Thanks for your help Dalvorn but there is no 3a only a 1/3a" you will multiply by 3a to get rid of your 1/3 a

OpenStudy (anonymous):

The idea of multiplying everything by 3a is to get rid of that fraction, since \[\frac{1}{3a}\times 3a=1\]

OpenStudy (anonymous):

The answer is -21 FYI

OpenStudy (anonymous):

can you rewrite the equation with the equation tool??

OpenStudy (anonymous):

1+3a=-6 -1 -1 --------- 0+1/3a=-7 1/3a=-7 3*1/3a=-7*3 cross out the first 3 and the 3 in 1/3 and its a=-21